Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

n(n+1)(n+2) divisibility by 8

Expert replies
by vikz_316 » Wed Nov 26, 2008 10:10 pm
Hello, can anyone help me figure out this problem
If an integer n is to be chosen at random from the integers 1 to 96, inclusive, what is the probability that n(n + 1)(n + 2) will be divisible by 8?
A. 1/4
B. 3/8
C. 1/2
D. 5/8
E. 3/4

I am getting the answer as 3/8, but the answer given is 5/8.
Join the discussion
Source: — Problem Solving |

by gmataug08 » Wed Nov 26, 2008 10:29 pm
any 3 consecutive number set having two even numbers would be divisible by 8

eg, 4*5*6 , 12*13*14 , 26*27*28......

96/2 => gives 48 even numbers , hence we will get 48 sets

next , any 3 consecutive number sets having having 2 odd numbers & one 8 or 8 divisible in them would be divisible by 8

eg, 7*8*9 , 15*16*17 .........

96/8 = 12 => that gives 12 sets

total we have 60 sets => 60/96 = 5/8
Join the discussion

by vikz_316 » Wed Nov 26, 2008 10:50 pm
I think i figured it out, but if someone can explain an easier method, i will be extremely grateful.

Here is my method.

maximum value n can have = 96
Values that n,n+1,n+2 can have are
8,16,24,32.......96 = 12 numbers total
7,15,23,31.......95 = 12 numbers total
6,14,22,30.......94 = 12 numbers total
4,12,20,28.......92 = 12 numbers total
2,10,18,26.......90 = 12 numbers total

12*5 = 60 values of n that are possible for which n(n+1)(n+2) will be divisible by 8.
Hence probability = 60/96 = 5/8.
Is there an easier method?
Join the discussion

by vikz_316 » Wed Nov 26, 2008 10:53 pm
thanks a lot my friend, your method definitely is easier than mine. Same idea, but it just didn't strike me. Thanks!
Join the discussion

by gmataug08 » Wed Nov 26, 2008 10:57 pm
glad it helped. :)
Join the discussion

by linfongyu » Thu Mar 19, 2009 8:58 pm
wow. how anyone can solve this - under pressure, in around 2 minutes - is beyond me. to whoever that can, you deserve a cookie.
Join the discussion

Re: n(n+1)(n+2) divisibility by 8

by farooq » Wed Oct 28, 2009 12:30 am
vikz_316 wrote:Hello, can anyone help me figure out this problem
If an integer n is to be chosen at random from the integers 1 to 96, inclusive, what is the probability that n(n + 1)(n + 2) will be divisible by 8?
A. 1/4
B. 3/8
C. 1/2
D. 5/8
E. 3/4

I am getting the answer as 3/8, but the answer given is 5/8.
In this questions the probability is results/total outcomes.

n(n+1)(n+2) is a multiple of 8 if,
1. n is a multiple of 8.
2. n+1 is a multiple of 8.
3. n+2 is a multiple of 8.
4. n(n+1) is a multiple of 8.
5. n(n+2) is a multiple of 8.
6. (n+1)(n+2). Ignore this case because we any of them make it a multiple of 8 and we already considered it above.

so from 1...96, we have total 12 numbers that are divisible by 8 or that are multiple of 8.

we have 5 cases, as mentioned above.
therefore total possibilities that n(n+1)(n+2) is

5*12/total experiments.

60/96 = 10/16 =5/8.
Regards,
Farooq Farooqui.
London. UK

It is your Attitude, not your Aptitude, that determines your Altitude.
Join the discussion

by ssmiles08 » Wed Oct 28, 2009 4:31 am
good posts. You can also do this problem by odds and evens.

If n is even, then the product is sure to be divisible by 8 (or three 2's) n is even so it is divisible by one 2. (n+2) is divisible by 4 since it has two 2's.

1/2 of all numbers are even.

If n is odd, (n+1) should be divisible by 8 for the product to be divisible. 96/8 = 12 possible numbers (n+1)

so 12/96 = 1/8 of all numbers are divisible by 8 when n is odd.

so you add both possibilities; when n is even and odd: 1/2 + 1/8 = 5/8
You got a dream... You gotta protect it. People can't do somethin' themselves, they wanna tell you you can't do it. If you want somethin', go get it. Period.
Join the discussion

by Mom4MBA » Fri Jan 22, 2010 1:18 pm
This is long method but easy to understand:

2,3,4
3,4,5
4,5,6
5,6,7
6,7,8
7,8,9 .....................n+1 multiple of 8
8,9,10
9,10,11
10,11,12
11,12,13
12,13,14
13,14,15
14,15,16
15,16,17 ................n+1 multiple of 8
.
.
.
96,97,98

here where n=even; n(n+1)(n+2) is divisible by 8, so we have 96/2 cases

Also see where n+1 is a multiple of 8, the set is divisible by 8, So between 2 - 97 (this is the range for n+1) we have 12 multiples of 8, thus we have 12 more cases

in all 96/2 + 12 = 48+12 = 60 cases are favorable
and total cases are 96

probability = 60/96 = 5/8 Answer
Join the discussion

by asherman » Fri Jan 22, 2010 4:09 pm
Thanks for all the info - it helped me a lot:)
Join the discussion